# HOPSE: Scalable Higher-Order Positional and Structural Encoder for Combinatorial Representations

### v1

@@ B.4 Beyond minimal routes: including adjacencies
Appendix C Proofs
Before proving Theorem 4.1, the following paragraphs show that HOPSE is at least as powerful as the WL test. Although this is a weaker result than the one from Theorem 4.1, the proof is more intuitive, making it a valuable tool to understand the HOPSE model.
First, a beneficial result from Gurugubelli and Chepuri [20] regarding the WL test is reported.
Lemma C.1.
For a graph ${\mathcal{G}}=({\mathcal{V}},{\mathcal{E}})$, the update of WL coloring of a node $\sigma\in{\mathcal{V}}$ at iteration K can be represented by the following equation
$c_{\sigma}^{(K)}=\phi\left(c_{\sigma}^{(0)},\{\mskip-5.0mu\{c_{\tau}^{(0)},\forall\tau\in\{{\mathcal{U}}^{(1)}({\sigma}),...,{\mathcal{U}}^{(K)}({\sigma})\}\}\mskip-5.0mu\}\right),$ (9)
where $\phi(\cdot)$ is some injective function and ${\mathcal{U}}^{(k)}(\sigma)$ denotes the set of upper adjacent $t$-hop neighbors of node $\sigma$.
Using this lemma, the following theorem can now be tackled.
Theorem C.2.
HOPSE is at least as powerful as the WL test in distinguishing non-isomorphic graphs if the functions $\varepsilon_{r}(\cdot)$, $g_{k}(\cdot)$, $\psi_{k}(\cdot)$, $f_{r,k}(\cdot)$, and $\Theta_{r}(\cdot)$ are injective for all $k\in{1,...,K}$ and for all $r\in{0,...,R}$.
Proof.
Let ${\mathcal{T}}_{1}$ and ${\mathcal{T}}_{2}$ be two graphs. By considering their nodes as zero rank cells and their edges as rank-1 cells, they can be considered as CC without loss of generalization. Given two nodes (zero rank cells) $\sigma_{1}\in{\mathcal{T}}_{1}$ and $\sigma_{2}\in{\mathcal{T}}_{2}$ let $c_{\sigma_{1}}^{(K)}$, $c_{\sigma_{2}}^{(K)}$ be the WL coloring of the two nodes after $K$ steps of the WL algorithm, and $h_{\sigma_{1}}^{(K)}$ and $h_{\sigma_{2}}^{(K)}$ be the embeddings obtained with HOPSE using a set of neighborhoods $\{{\mathcal{N}}_{1},\dots,{\mathcal{N}}_{T}\}$. Showing that for all graphs ${\mathcal{T}}_{1}$ and ${\mathcal{T}}_{2}$ and for all nodes $\sigma_{1}$ and $\sigma_{2}$, there exists a version of HOPSE such that $h_{\sigma_{1}}^{(K)}=h_{\sigma_{2}}^{(K)}$ implies $c_{\sigma_{1}}^{(K)}=c_{\sigma_{2}}^{(K)}$ proves that HOPSE is at least as powerful as the WL isomorphism test. This is because it would mean that the WL test fails to assign different embeddings whenever HOPSE fails.
As the focus is solely on the embeddings of the nodes in the considered graphs, the target rank is fixed to $r_{k}=0$, which is omitted in the remainder of the proof for clarity. Furthermore, since only neighborhoods ${\mathcal{N}}_{k}$ with target rank 0 are relevant, the proof is restricted to the case ${\mathcal{N}}_{k}={\mathcal{U}}^{(k)}$ for $k\in{1,\dots,K}$, where ${\mathcal{U}}^{(k)}$ denotes the upper adjacent $k$-hop neighborhood.
Without loss of generality, the analysis can be restricted to strictly augmented multi-attributed Hasse graphs defined inSection 4 of the form ${\mathcal{J}}_{{\mathcal{N}}}=({\mathcal{C}}_{{\mathcal{N}}},{\mathcal{E}}_{{\mathcal{N}}},{\mathcal{F}}_{{\mathcal{N}}})$, where ${\mathcal{F}}_{{\mathcal{N}}}={\mathcal{F}}$ for all ${\mathcal{N}}\in N^{\rightarrow 0}$, and ${\mathcal{F}}$ denotes the initial node features of the graph. As only a single feature type is considered, the index $k=0$ is omitted in the remainder of the proof for clarity.
For any node $\sigma$, by combining equation 4 to equation 5 the HOPSE embeddings can be written as
$h_{\sigma}^{(K)}=\Theta\Big{(}\varepsilon(\mathbf{Z}),\{f\big{(}\psi(\{g(\sigma,{\mathcal{J}}_{{\mathcal{N}}_{j}})|{\mathcal{N}}_{j}\in N^{\rightarrow 0}\})\big{)}\}\Big{)}.$ (10)
Since $g({\mathcal{J}}_{{\mathcal{N}}})$ is the function that creates the encodings for all target cells in the graph by creating a row for each cell, $g(\sigma,{\mathcal{J}}_{{\mathcal{N}}})$ is defined as the row-wise version of $g({\mathcal{J}}_{{\mathcal{N}}})$.
The functions $g(\sigma,{\mathcal{J}}_{{\mathcal{N}}})$ can be any injective functions, but by considering only functions that depend on the 1-hop neighbors of $\sigma$ in the corresponding Hasse graph, it is immediate to see that $g(\sigma,{\mathcal{J}}_{{\mathcal{U}}_{j}})=\xi(\{\mskip-5.0mu\{{\mathcal{F}}[\tau]|\forall\tau\in{\mathcal{U}}^{(j)}(\sigma)\}\mskip-5.0mu\})$, given that ${\mathcal{N}}_{j}={\mathcal{U}}^{(j)}$. Since by hypothesis $g(\cdot)$ is injective, also $\xi(\cdot)$ is. Since $\varepsilon(\cdot)$ is also injective by hypothesis equation 10 can be rewritten as
$\begin{split}h_{\sigma}^{(K)}&=\hat{\Theta}\left(c_{\sigma}^{(0)},\{\mskip-5.0mu\{c_{\tau}^{(0)},\forall\tau\in{\mathcal{U}}^{(1)}(\sigma)\}\mskip-5.0mu\},...,\{\mskip-5.0mu\{c_{\tau}^{(0)},\forall\tau\in{\mathcal{U}}^{(K)}(\sigma)\}\mskip-5.0mu\}\right)=\\ &=\hat{\Theta}\left(c_{\sigma}^{(0)},\{\mskip-5.0mu\{c_{\tau}^{(0)},\forall\tau\in\{{\mathcal{U}}^{(1)}(\sigma),...,{\mathcal{U}}^{(K)}(\sigma)\}\}\mskip-5.0mu\}\right),\end{split}$ (11)
where the fact that $c_{\sigma}^{(0)}=\mathbf{Z}$ was used, which means that for the WL coloring algorithm the nodes’ features are the initial features. Also, since the functions $\Theta(\cdot)$, $f(\cdot)$, $\psi(\cdot)$, and $\xi(\cdot)$ are injective, $\hat{\Theta}(\cdot)$ is also injective since it is a composition of injective functions.
Since $\hat{\Theta}$ is injective $h_{\sigma_{1}}^{(K)}=h_{\sigma_{2}}^{(K)}$ implies that the arguments of $\hat{\Theta}$ are equal. But those arguments are the exact arguments for the WL colorings of nodes $\sigma_{1}$ and $\sigma_{2}$ reported in equation 9, which means that $h_{\sigma_{1}}^{(K)}=h_{\sigma_{2}}^{(K)}$ implies $c_{\sigma_{1}}^{(K)}=c_{\sigma_{2}}^{(K)}$. As discussed previously, this implies that HOPSE is at least as powerful as the WL test. ∎
The definition of the CCWL test, which generalizes the WL test to combinatorial complexes, is taken from [39].
Definition C.3 (The CC Weisfeiler-Leman (CCWL) test on labeled combinatorial complexes).
Let $({\mathcal{T}},F)$ be a labelled combinatorial complex. Let ${\mathcal{N}}$ be a neighborhood on ${\mathcal{C}}$. The scheme proceeds as follows:
  • Initialization: Cells $\sigma$ are initialized with the labels given by F, i.e.: for all $\sigma\in{\mathcal{C}}$: $c_{\sigma,F}^{(0)}=F(\sigma)$.
  • Refinement: Given colors of cells at iteration $k$, the refinement step computes the color of cell $\sigma$ at the next iteration $c_{\sigma,F}^{(k+1)}$ using a perfect HASH function as follows:
    $\begin{split}c_{\mathcal{N}}^{(k)}(\sigma)&=\{\mskip-5.0mu\{c_{\tau,F}^{(k)}|\forall\tau\in{\mathcal{N}}(\sigma)\}\mskip-5.0mu\},\\ c_{\sigma,F}^{(k+1)}&=\text{HASH}\left(c_{\sigma,F}^{(k)},c_{\mathcal{N}}^{(k)}(\sigma)\right).\end{split}$ (12)
  • Termination: The algorithm stops when an iteration leaves the coloring unchanged.
To show that HOPSE is more powerful than the CCWL test, the following lemma is proven first.
Lemma C.4.
For a combinatorial complex $({\mathcal{T}},F)$ the update of CCWL coloring of a cell $\sigma\in{\mathcal{C}}$ at iteration $T$ can be represented as the following equation
$c_{\sigma}^{(K)}=\phi\left(c_{\sigma}^{(0)},\{\mskip-5.0mu\{c_{\tau}^{(0)},\forall\tau\in\{{\mathcal{N}}^{(1)}({\sigma}),...,{\mathcal{N}}^{(K)}({\sigma})\}\}\mskip-5.0mu\}\right),$ (13)
where $\phi(\cdot)$ is some injective function, and ${\mathcal{N}}^{(k)}$ denotes the $t$-hops version of the neighborhood ${\mathcal{N}}$.
Proof.
The proof is by induction. For $k=0$, the equation holds trivially, since the initial coloring for cell $\sigma$ is $c_{\sigma}^{(0)}$. Assuming that the $k-1$ update is true
$c_{\sigma}^{(t-1)}=\phi_{0}\left(c_{\sigma}^{(0)},\{\mskip-5.0mu\{c_{\tau}^{(0)},\forall\tau\in\{{\mathcal{N}}^{(1)}(\sigma),...,{\mathcal{N}}^{(t-1)}(\sigma)\}\}\mskip-5.0mu\}\right).$
By substituting this in equation 12 the $k$ update can be written as
$\begin{split}c_{\sigma}^{(k)}=\text{HASH}\Big{(}&\phi_{0}\big{(}c_{\sigma}^{(0)},\{\mskip-5.0mu\{c_{\tau}^{(0)},\forall\tau\in\{{\mathcal{N}}^{(1)}(\sigma),...,{\mathcal{N}}^{(t-1)}(\sigma)\}\}\mskip-5.0mu\}\big{)},\\ \{\mskip-5.0mu\{&\phi_{0}\big{(}c_{\upsilon}^{(0)},\{\mskip-5.0mu\{c_{\tau}^{(0)},\forall\tau\in\{{\mathcal{N}}^{(1)}(\sigma),...,{\mathcal{N}}^{(t-1)}(\sigma)\}\}\mskip-5.0mu\}\big{)},\forall\upsilon\in{\mathcal{N}}^{(1)}(\sigma)\}\mskip-5.0mu\}\Big{)}.\end{split}$
Now, since HASH and $\phi_{0}$ are injective functions, it follows that
$\begin{split}c_{\sigma}^{(k)}&=\phi\Big{(}c_{\sigma}^{(0)},\{\mskip-5.0mu\{c_{\tau}^{(0)},\forall\tau\in\{{\mathcal{N}}^{(1)}(\sigma),...,{\mathcal{N}}^{(t-1)}(\sigma)\}\}\mskip-5.0mu\},\\ &\hskip 14.22636pt\{\mskip-5.0mu\{\big{(}c_{\upsilon}^{(0)},\{\mskip-5.0mu\{c_{\tau}^{(0)},\forall\tau 1\in\{{\mathcal{N}}^{(1)}(\sigma),...,{\mathcal{N}}^{(t-1)}(\sigma)\}\}\mskip-5.0mu\}\big{)},\forall\upsilon\in{\mathcal{N}}^{(1)}(\sigma)\}\mskip-5.0mu\}\Big{)}\\ &=\phi\Big{(}c_{\sigma}^{(0)},\{\mskip-5.0mu\{c_{\tau}^{(0)},\forall\tau\in\{{\mathcal{N}}^{(1)}(\sigma),...,{\mathcal{N}}^{(t-1)}(\sigma)\}\}\mskip-5.0mu\},\{\mskip-5.0mu\{c_{\tau}^{(0)},\forall\tau\in\{{\mathcal{N}}^{(1)}(\sigma),...,{\mathcal{N}}^{(k)}(\sigma)\}\}\mskip-5.0mu\}\Big{)}\\ &=\phi\Big{(}c_{\sigma}^{(0)},\{\mskip-5.0mu\{c_{\tau}^{(0)},\forall\tau\in\{{\mathcal{N}}^{(1)}(\sigma),...,{\mathcal{N}}^{(k)}(\sigma)\}\}\mskip-5.0mu\}\Big{)}\end{split}$
This is the result needed, so the theorem is proven by induction. ∎
This result is the combinatorial complex equivalent of Lemma C.1, and it can now be used to prove the following theorem.
Theorem C.5.
HOPSE is at least as powerful as the CCWL test in distinguishing non-isomorphic combinatorial complexes if the functions $\varepsilon_{r}(\cdot)$, $g_{k}(\cdot)$, $\psi_{k}(\cdot)$, $f_{r,k}(\cdot)$, and $\Theta_{r}(\cdot)$ are injective for all $k\in{1,...,K}$ and for all $i\in{0,...,R}$.
Proof.
Given that Lemma C.4 is equivalent to Lemma C.1 with ${\mathcal{N}}_{i}={\mathcal{N}}^{(i)}$ for $i\in 0,...,K$, this proof follows the same steps as the proof for Theorem C.2. ∎
Figure 4: a. Two combinatorial complexes with nodes, edges, and faces. ${\mathcal{T}}_{1}$ is a icosahedron while ${\mathcal{T}}_{2}$ is composed of five tetrahedrons. b. Their respective Hasse graphs given the choice of the neighborhood ${\mathcal{A}}_{2,1}$, which connects faces that share an edge. Image adapted from [39].
Now the two main theorems of the paper can be proven.
Theorem (Theorem 4.1).
HOPSE is more powerful than the CCWL test in distinguishing non-isomorphic combinatorial complexes if the functions $\varepsilon_{r}(\cdot)$, $g_{k}(\cdot)$, $\psi_{k}(\cdot)$, $f_{r,k}(\cdot)$, and $\Theta_{r}(\cdot)$ are injective for all $k\in{1,...,K}$ and for all $r\in{0,...,R}$.
Proof.
Since HOPSE matches the expressive power of CCWL, as established in Theorem C.5, to prove this theorem is sufficient to present a case in which two combinatorial complexes are indistinguishable by the CCWL test but are successfully distinguished by the HOPSE algorithm. The example in Figure 4 illustrates two such CCs. Using the neighborhood ${\mathcal{A}}_{2,1}$ (the faces are adjacent if they share an edge), the corresponding Hasse graphs ${\mathcal{J}}_{1}$ and ${\mathcal{J}}_{2}$ are constructed, both of which are 3-regular. It is known that the WL test cannot distinguish between regular graphs of the same degree [29]. According to Proposition B.11 in [39], the CCWL test on these CCs is equivalent to the WL test on ${\mathcal{J}}_{1}$ and ${\mathcal{J}}_{2}$, and thus also fails to distinguish them. However, the HOPSE algorithm can separate the two cases. In fact, by choosing the function $g_{1}(\cdot)$ to return the size of the connected component to which a node belongs, the difference between the two HOPSE encodings is apparent. ∎
Theorem (Theorem 4.2).
HOPSE is cell permutation equivariant if the functions $\varepsilon_{r}(\cdot)$, $g_{k}(\cdot)$, $\psi_{k}(\cdot)$, $f_{r,k}(\cdot)$, and $\Theta_{r}(\cdot)$ are cell permutation equivariant for all $k\in{1,\dots,K}$ and all $r\in{0,\dots,R}$.
Proof.
To demonstrate that HOPSE is cell permutation equivariant, it suffices to show that the composition of permutation equivariant functions remains permutation equivariant. Since, by assumption, all functions that compose HOPSE are permutation equivariant, the result follows directly. Let $\pi(\cdot)$ be the permutation operator. Then $p(\cdot)$ is permutation equivariant if $\pi(p(\cdot))=p(\pi(\cdot))$. Given two permutaiton equivariant functions $p(\cdot)$ and $q(\cdot)$ their composition $r(\cdot)=(p\circ q)(\cdot)$ is also permutation equivariant:
$r(\pi(\cdot))=(p\circ q)(\pi(\cdot))=p(q(\pi(\cdot)))=p(\pi(q(\cdot)))=\pi(p(q(\cdot)))=\pi(r(\cdot)).$
@@ Appendix D Framework realization
This section contains an in-depth description of the algorithmic procedure for extracting structural and positional information according two the proposed realizations HOPSE-M and HOPSE-G. Complementary to Section 4.3 the following is an in-depth explanation of the procedure for computing the positional and structural encodings of both realizations.
The first step in the pipeline is the definition of the family of functions $(g_{0},\dots,g_{K})$ and the same target aggregations $\psi_{k}(\cdot)$. Both the variations have a different way of calculating the output embeddings so both of the cases are shown.
D.1 Precomputation of HOPSE-M
In this case four hand-picked structural and positional encodings were used. Let ${\mathcal{G}}=(V,E)$ be an undirected graph with $|V|=n$ nodes. The individual channels $g_{k}({\mathcal{G}})$ are defined as follows:
@@ D.1 Precomputation of HOPSE-M
    $g_{1}(\cdot)=LapPE_{i}=[\phi_{1},\dots,\phi_{i}]\in\mathbb{R}^{n\times i},$ the first $i$ nontrivial eigenvectors of the graph Laplacian (up to frequency $i$).
@@
    $g_{2}(\cdot)=RWSE_{K}\in\mathbb{R}^{n\times K},$ the Random Walk Structural Encodings, obtained by stacking the diagonal entries of the $t$-step random-walk transition matrices for $t=1,\dots,K$ (here $K=60$).
@@
    $g_{3}(\cdot)=HKdiagSE_{K}\in\mathbb{R}^{n\times K},$ the Heat Kernel Diagonal Structural Encodings, given by $[\exp(-\lambda_{j}\,t)]_{j=1,\dots,n;\,t=1,\dots,K}$ for the Laplacian eigenvalues $\{\lambda_{j}\}$.
@@
    $g_{4}(\cdot)=ElstaticPE\in\mathbb{R}^{n\times d_{4}},$ the Electrostatic Potential Positional Encodings, computed by solving the discrete Poisson equation on ${\mathcal{G}}$ with suitable boundary conditions.
Importantly, since each $g_{k}$ depends only on the adjacency structure of ${\mathcal{G}}$, none of these channels requires access to any node feature vectors. For more indepth explanation of each PS encoding, see [13].
D.2 Precomputation of HOPSE-G
@@ D.3 Learnable embedding functions
All MLPs share depth $L$ but may vary in hidden-layer widths $\{d_{r,\ell}\}$. This unified design—with layer normalization before activation—ensures stable gradient flow and flexible expressivity across channels and ranks.
@@ Appendix E Complexity Analysis
To analyze the time complexity (in terms of FLOPs) of the HOPSE framework, distinguishing between two components is needed: (i) operations that can be precomputed once before training, and (ii) operations that must be performed during each training iteration.
@@
For positional and structural encodings, although the functions $g_{k}(\cdot)$ can, in principle, be any function—including trainable neural networks—the implementation is restricted to either graph-theoretic positional encodings or those generated by a pretrained GPSE model. This allows the tensors $\mathbf{X}_{r,k}$ (as defined in equation 2) to be pre-computed before training begins.
@@
The trainable components of the HOPSE framework are the functions $\varepsilon_{r}(\cdot)$, $f_{r,k}(\cdot)$, and $\Theta_{r}(\cdot)$, for all $r=1,\dots,R$ and $k=1,\dots,K$. For these functions, HOPSE uses MLPs with $L$ layers
An MLP with $L$ layers applied to $N$ elements with input and output dimensions $D_{\text{in}}$ and $D_{\text{out}}$ has a computational complexity of ${\mathcal{O}}(ND_{\text{in}}D_{\text{out}})$. For simplicity, let us consider fixed input and output dimensions for each rank. The complexity of generating the initial features $\hat{\mathbf{Z}}_{r}$ in equation 4 is
@@
Next, the cost of computing $\mathbf{X}_{r,k}$ for all $K$ positional encodings (as in equation 3) is:
@@
Assuming concatenation is used as the aggregation operation in equation 5, and that the output features have dimension $D_{r}$, the complexity of computing the final embeddings $\mathbf{H}_{r}$ is:
@@
This complexity scales linearly with the number of rank-$r$ cells and is independent of their connectivity. This is in sharp contrast with most higher-order neural architectures. For example, the Generalized Combinatorial Complex Neural Network (GCCN) analyzed in [39] has the following complexity:
@@
Summary.
This analysis shows that HOPSE achieves high expressive power (as demonstrated in Theorem 4.1) while maintaining a computational complexity that scales linearly with the number of cells. This is possible by offloading the structure-dependent computations to a one-time preprocessing step. The resulting efficiency, confirmed by the experimental results in Table 3, makes HOPSE well-suited for large-scale graph data.
Appendix F Extended results

### v2

@@ Appendix D Framework realization
This section contains an in-depth explanation of the empirical choices of PSEs that realize the theoretical results of this work, serving as a complement to Section˜4.2. It is important to recall that the previous theoretical results hold for any choice of functions $g_{k}(\cdot)$ and $\psi_{k}(\cdot)$ (characterization and aggregation, respectively), as long as they fulfill the injectivity and permutation equivariance constraints. The fact that a certain choice of functions does not completely fulfill the injectivity or permutation equivariance constraints does not imply that they have to be completely disregarded. These functions are still relevant to study the empirical reach of the theoretical results in a more constrained setting. More importantly, fully injective and permutation equivariant functions need to be learned (or fine-tuned) while training, greatly impacting the computational advantage of this method. As such, these endeavours are defered to future works.
D.1 Handcrafted PSEs
In the literature, functions over graphs that characterize their structure are commonly referred to as positional and structural encodings (PSEs). While it is true that not all PSEs are meant to act as fully characterizing functions, their purpose is to represent certain characteristics of the graph that are implicitly present in its structure. Constructing PSEs that fully characterize graphs up to isomorphism and preserve injectivity results in the graph canonization problem Dong et al. [2023]. Consequently, there will always be a trade-off between the goodness of the characterization and the computational complexity. With this in mind, the following choices of characterization functions stem from the literature as commonly used methods to augment the performance of GNNs. Additionally, while not fully injective, they are permutation equivariant, completely fulfilling that theoretical criterion.
Let ${\mathcal{G}}=(V,E)$ be an undirected graph with $|V|=n$ nodes. The characterization functions $g_{k}:{\mathcal{G}}\to\mathbb{R}^{d\times d}$ with arbitrary codomains used in this work are defined as follows:
@@ D.1 Handcrafted PSEs
    $g_{1}(\cdot)=\mathrm{LapPE}_{i}=[\phi_{1},\dots,\phi_{i}]\in\mathbb{R}^{n\times i},$ the first $i$ nontrivial eigenvectors of the graph Laplacian (up to frequency $i$).
@@
    $g_{2}(\cdot)=\mathrm{RWSE}_{K}\in\mathbb{R}^{n\times K},$ the Random Walk Structural Encodings, obtained by stacking the diagonal entries of the tt‑step random‑walk transition matrices for $t=1,\dots,K$ (here $K=60$).
@@
    $g_{3}(\cdot)=\mathrm{HKdiagSE}_{K}\in\mathbb{R}^{n\times K},$ the Heat Kernel Diagonal Structural Encodings, given by $\exp(-\lambda_{j}\,t)]_{j=1,\dots,n;\,t=1,\dots,K}$ for the Laplacian eigenvalues $\{\lambda_{j}\}$.
@@
    $g_{4}(\cdot)=\mathrm{ElstaticPE}\in\mathbb{R}^{n\times d},$ the Electrostatic Potential Positional Encodings, computed by solving the discrete Poisson equation on $\mathcal{G}$} with suitable boundary conditions.
Importantly, since each $g_{k}(\cdot)$ depends only on the adjacency structure of $\mathcal{G}$}, none of these channels requires access to any node feature vectors. For more indepth explanation of each PSE, see Cantürk et al. [2023].
D.2 Pretrained encodings
@@ D.3 Learnable embedding functions
All MLPs share depth $L$ but may vary in hidden‐layer widths $\{d_{r,\ell}\}$. This unified design—with layer normalization before activation—ensures stable gradient flow and flexible expressivity across channels and ranks.
@@ Appendix E Complexity analysis
The computational complexity of HOPSE scales linearly with the number of cells. This is possible by offloading the structure-dependent computations to a one-time preprocessing step. The resulting efficiency, confirmed by the experimental results in Table˜4, makes HOPSE well-suited for large-scale graph data. The following sections break down the process into two distinct parts (i.) preprocessing and (ii.) training time complexity.
@@
For positional and structural encodings, although the functions $g_{k}(\cdot)$ can, in principle, be any function (including trainable neural networks) the implementation is restricted to either graph-theoretic positional encodings or those generated by a pretrained GPSE model. This allows the tensors $\mathbf{X}_{r,k}$ (as defined in equation 2) to be pre-computed before training begins.
@@
The trainable components of the HOPSE framework are the functions $\varepsilon_{r}(\cdot)$, $f_{r,k}(\cdot)$, and $\Theta_{r}(\cdot)$, for all $r=1,\dots,R$ and $k=1,\dots,K$. HOPSE M and G use MLPs with $L$ layers to model these functions.
These MLPs applied to $N$ elements with input and output dimensions $D_{\text{in}}$ and $D_{\text{out}}$ have a computational complexity of ${\mathcal{O}}(ND_{\text{in}}D_{\text{out}})$. For simplicity, let us consider fixed input and output dimensions for each rank. The complexity of generating the initial features $\hat{\mathbf{Z}}_{r}$ in Equation˜4 is
@@
Next, the cost of computing $\mathbf{X}_{r,k}$ for all $K$ positional encodings (as in Equation˜3) is:
@@
Assuming concatenation is used as the aggregation operation in Equation˜5, and that the output features have dimension $D_{r}$, the complexity of computing the final embeddings $\mathbf{H}_{r}$ is:
@@
This complexity scales linearly with the number of rank-$r$ cells and is independent of their connectivity. This is in sharp contrast with most higher-order neural architectures. For example, the Generalized Combinatorial Complex Neural Network (GCCN) analyzed in Papillon et al. [2024] has the following complexity:
@@
Appendix F Experimental configuration